Marcus delivers papers for the local newspaper. Usually he can deliver all of the papers in 90 minutes. However, today he slept in and is running late. Marcus’s dad helps him and can deliver all of the papers in 2 hours. If Marcus and his dad only have 1 hour to deliver the papers, will they deliver them all in time?
Answer by jorel1380(2518)

n/1.5+n/2=1
4n+3n=6
7n=6
n=6/7 hours to deliver all the papers
Yes, they will have time to deliver all the papers..